OpenStudy (anonymous):

Approximate further to a linear form at x~0, I am lost at this step

OpenStudy (anonymous):

\[y(x) \approx \frac {x} {x(1+x)}\]

OpenStudy (anonymous):

\[\approx \frac {x}{x(1+x)} + \frac {x(x-2)}{2(x+x^2)^2} \times (X-x) \approx {x \over x}+\frac {x(-2)}{2x^4}\times (X-x) \approx 1 - X\]

OpenStudy (anonymous):

Is this how its supposed to be done?

OpenStudy (unklerhaukus):

\[y(x)=\frac{x}{x(1+x)}=\frac{1}{1+x}\]

OpenStudy (anonymous):

= 1 - x so simple, can't believe I spend an hour on it :(

OpenStudy (unklerhaukus):